# Evolutionary Computation in the Era of Large Language Model: Survey and Roadmap

### v1

@@ IV-D Other Generative Tasks
Some research focuses on text generation. Xiao et al. used the generation ability of LLM and the search ability of EA for news summary generation [35]. They utilized event patterns structure to represent and describe the key information, relationships and characteristics of events. This study used LLM to extract event patterns from text, and then used genetic algorithm to evolve the event pattern pool according to the importance and information quantity of argument roles, selecting the event pattern with the highest fitness to input into LLM for news summary generation. A similar application is to generate Super Mario game levels using LLMs [101, 32]. Sudhakaran et al. used a string representation method similar to the Video Game Level Corpus dataset [102] to encode levels, and then trained a LLM namely MarioGPT based on GPT-2 that can generate Mario levels according to natural language descriptions. By combining MarioGPT with the Novelty Search algorithm [103], an open-ended level generation process was formed. Novelty search can continually discover levels with diverse structures and gameplay styles.
Berger et al. focused on the text-to-image generation task. In this type of task, when using LLMs like Stable Diffusion [104] to generate photorealistic images, it is necessary to optimize the input text prompts and model parameters in order to achieve better performance. Berger et al. tried to solve this problem and proposed the StableYolo method [105]. Its main innovation lies in combining EAs with computer vision for the first time, simultaneously optimizing language prompts and model hyperparameters to improve image generation quality. Among them, the LLM Stable Diffusion plays the main role of image generation, generating a series of images based on the input text prompts. While the Genetic Algorithm plays the role of searching and optimizing. It uses the recognition confidence given by the YOLO object detection model as the fitness function to perform multi-objective optimization. The optimization objects include keywords in text prompts and various hyperparameters of the Stable Diffusion model. Through repeated iteration, the best language prompts and model settings that maximize image quality are found.
Different from aforementioned studies, Suzuki et al. applied LLMs and EAs to study the evolutionary dynamics of social populations under conflicts of interest and behaviors between individuals [106], i.e. the propagation and changes of different personality traits and behavioral strategies. Suzuki et al. used LLMs to make trait expressions more complex and higher-order, able to directly describe personality and psychological attributes that are difficult to model directly, and map them to behaviors. They then constructed an agent-based model where each agent uses a short natural language description to represent a personality trait related to cooperation as its gene. The LLM was used to derive a behavioral strategy for each agent based on the personality trait description and game history. The population evolved according to a genetic algorithm, with offspring inheriting mutated versions of their parents’ trait genes, thus simulating the evolutionary dynamics of population changes over generations. This study provided a novel approach for researching the evolution of complex social populations.
@@ V Roadmap and Future Directions
In the previous sections, we have reviewed the recent advances in unifying LLMs and EAs, but there are still many challenges and open problems that need to be addressed. In this section, we discuss the future directions of this research area.
V-A LLM-assists EA for Complex and Large-scale Optimization
@@
Existing research has partially validated the ability of LLMs to address small-scale numerical optimization problems [13, 14, 15], but practicality still poses challenges. Firstly, for complex optimization problems with high-dimensional search spaces, numerous constraints, and high-precision numerical optimization, the limited context understanding and prompt length restrictions of large models may increase the difficulty of interaction with LLMs. Secondly, as black-box optimizers, the decision-making process of LLMs is difficult to interpret [107], and it remains unknown whether their optimization ability stems from reasoning or randomness. Moreover, the scope of the evaluated problems in existing research is relatively narrow [13, 14, 15], often limited to specific optimization problems, and the assessments only consider limited influencing factors, which are insufficient to fully demonstrate the optimization capabilities of LLMs. Another issue is that empirical studies suggest that LLMs may struggle to handle constrained problems effectively, making it crucial to incorporate constraint conditions into LLM optimization frameworks.
@@
Therefore, to further explore and utilize the capabilities of LLMs in solving optimization problems, future work can focus on addressing the aforementioned challenges and improving various aspects such as LLM behavior interpretation and evaluation, rational utilization of optimization information, and solving complex problems. Currently, comprehensive, fair, and adequately diverse evaluations are needed to assess the optimization capabilities of different LLMs on a wide range of optimization problems. Additionally, efforts can be made to interpret and understand the behavior of LLMs, such as analyzing their internal attention mechanisms [108] or investigating the relationship between LLM temperature settings and exploration of optimal solutions, in order to gain a more objective understanding of LLM optimization abilities. Based on these foundations, improvement strategies or new pre-training models can be proposed to solve optimization problems with higher dimensions and complex constraint conditions. Beyond these research directions, the contextual understanding and processing abilities of LLMs can be further exploited, for example, by analyzing historical information of the optimization process and considering more advanced prompt methods [109].
In the task of EA generation based on LLMs, similar challenges arise when LLMs are involved in the analysis and understanding of optimization problems. Existing LLMs may struggle to handle complex problem descriptions and large numerical inputs effectively. Furthermore, the analytical capabilities of LLMs are limited by their own proficiency, which can impact the quality of modeling and encoding [36]. In addition, there are challenges in the generation and debugging stages of algorithm development using LLMs. For instance, interaction-based approaches for automated algorithm design with LLMs may suffer from low efficiency, as each interaction can be time-consuming [37]. Moreover, LLM-based debugging is limited to fixing minor errors and cannot address fundamental algorithmic flaws, such as model errors [37]. To overcome these challenges, future research can explore the development of new pre-training models specifically designed to address these issues. These models can enhance the LLMs’ handling of complex problem descriptions, improve their ability to process large numerical inputs, and enhance their analytical capabilities for more accurate modeling and encoding. Additionally, efforts can be made to improve the efficiency of LLM-based interactions in algorithm design, and explore methods to address fundamental algorithmic errors through LLM-based debugging, which will be discussed in Section V-F in details.
@@ V-B LLM-assisted Automatous and Intelligent EA
In addition to the problems mentioned above, applying LLMs to enhance evolutionary algorithms can bring broader advantages.
First, multi-modal LLMs provide opportunities for cross-domain EAs. Some LLMs have learned the relationships between different modalities during pre-training, allowing them to be used for downstream tasks involving multiple modalities [110, 111], such as image captioning and text-to-image generation. The ability of LLMs to understand correspondences across modalities can help evolutionary algorithms achieve cross-modal crossover and mutation by providing cross-modal prompts or examples. Additionally, population construction and evaluation can be improved across modalities. Furthermore, as LLMs technologies continue to progress, evolutionary operators based on LLMs will also benefit and improve the performance of EAs. With increasing model scales and stronger generalization capabilities, evolutionary operators based on LLMs will be better able to simulate more complex evolutionary mechanisms precisely. This provides potential for tackling problems in complex search spaces. The emerging and rapidly advancing field of LLMs will also provide more choices for EAs. LLMs with different pre-training characteristics will continue to enrich and enhance the capabilities of evolutionary operators based on LLMs. More complex optimization problems such as multi-objective optimization [112], transfer optimization [113], multi-task optimization [114], and multiform optimization [115] may achieve new performance improvements.
@@
V-C Theoretical Study and Empirical Analysis for Specific LLM+EA Methods
Although empirical studies have validated the effectiveness of combining LLM and EA on small-scale problems, the incentives of their interaction are not yet clear. This reminds us to explore the sources of mutual promotion between LLM and EA in theoretical research, and to analyze in detail their complementary advantages and existing issues in large-scale empirical studies, thereby further promoting algorithm performance improvement. This research can delve into the following aspects:
V-C1 Algorithm Analysis
Researchers can analyze specific algorithms that combine LLM and EA and explore their convergence, complexity, and other properties. For convergence analysis, researchers can verify whether the algorithm can converge to the optimal solution or local optimum during the iterative process. This may involve proving the optimization properties of the objective function or fitness function during the algorithm’s iterative process and analyzing the impact of algorithm parameter settings on convergence. Regarding complexity analysis, future work can analyze the time and space complexity of algorithms that combine LLM and EA. This includes evaluating the complexity of key steps in the algorithm, such as generating new individuals, selection, crossover, and mutation operations.
V-C2 Optimization Theory
In problem modeling and analysis, specific problems that combine LLM and EA can be modeled and analyzed from the perspective of optimization theory. This may involve defining and characterizing the objective function, constraints, and exploring the feasible solution space of the problem. The key point of research lies in the search strategy, where researchers can study how to select and design appropriate search strategies to improve the performance of LLM and EA combination. This may include comparing and analyzing the effects of different search strategies (such as selection, crossover, and mutation operations) and sensitivity analysis of search strategy parameters.
V-C3 Theoretical Limits
Firstly, the applicability of methods that combine LLM and EA on different types of problems can be studied, which may involve analyzing the relationship between problem attributes, constraints, problem size, and algorithm performance, as well as studying the performance guarantees or theoretical limits of algorithms on specific problem types. Moreover, future work can analyze the complexity of these methods on specific problems, including analyzing the computational complexity category of problems (such as P, NP, NP-hard) and analyzing the approximate performance of algorithms to determine the theoretical limits of algorithms in solving specific problems.
V-C4 Problem Characterization Study
Attribute analysis focus on how different problem attributes affect the performance of combining LLM and EA, which involves analyzing the characteristics, structure, and scale of problems, as well as how to select appropriate algorithms or adjust algorithm parameters based on problem attributes. Second is comparative research. Researchers can compare the performance of these methods with other optimization methods on different problems, which can be done by experimentally comparing the performance of different algorithms in solving specific problems, thereby gaining a deeper understanding of the advantages and limitations of combining LLM and EA methods.
By conducting in-depth research on the above aspects from both theoretical and experimental perspectives, we can comprehensively understand the research combining LLM and EA and promote the theoretical development and application exploration in this field.
V-D Robust and Stable Prompt Engineering
@@
To further enhance the performance of LLM and harness EAs to advance prompt engineering, future work can focus on addressing the aforementioned challenges to develop more effective methods. Regarding the issue of initialization, future research can consider multi-source initialization, simultaneously utilizing LLM to automatically expand the size and quality of the initial population. For problems with a high search space complexity, it is necessary to design more efficient evolutionary strategies, such as introducing richer evolutionary operators, leveraging the strengths of different EAs, and employing adaptive evolution. Another promising direction worth exploring is the utilization of human feedback to expedite and optimize the prompt discovery process, which has been overlooked in current research. In addressing the constraint of LLM on prompt performance, apart from considering the use of more powerful LLM models or designing dedicated fine-tuning methods, an alternative research direction at a lower level is to extract internal representations of LLM and study their mechanisms for understanding prompts, thereby guiding the design and representation of prompt strategies more effectively. Stable prompt optimization is also a promising direction to mitigate performance fluctuations caused by randomness.
V-E Efficient and Effective NAS
In the realm propelled by the collaboration of LLM and EA, NAS stands out as a vital application scenario. The prior knowledge about network architecture and code generation capability of LLM can greatly assist EA in efficiently discovering optimal architectures. However, existing work still faces various challenges. Some of these challenges are common among NAS methods based on EAs, such as high time consumption [116]. Additionally, the integration of LLM introduces new issues. Firstly, it is important to note that the current state-of-the-art LLM models, while excelling in various tasks, are not specifically tailored for NAS. Their ability to tackle NAS tasks stems from the presence of network architecture-related information in the training data [71]. Consequently, different LLM models exhibit significant variations in performance in NAS tasks. Furthermore, when compared to mainstream NAS methods, LLM-based approaches still exhibit some gaps in terms of their application scope and generalization ability. Additionally, certain ablation experiments reveal that the direct utilization of LLM prompts yields suboptimal learning results [76], underscoring the limitations of the current capabilities of LLM. Exploring fine-tuning methods that combine LLM with evolutionary algorithms has proven necessary.
V-F Generalized Code Generation for Complex Task
@@
LLM and EA jointly contribute to another field of advancement, i.e., code generation. Currently, a plethora of research has emerged in this domain, which has further spurred the development of various downstream tasks, such as software engineering and evolutionary algorithm design.
@@ VI Conclusion
In this paper, we embarked on a comprehensive exploration of the intersection between EAs and LLMs within the transformative era marked by advanced models like GPT-4. Our study identified key advantages of EAs, such as flexibility, global search capabilities, and iterative optimization mechanisms, which complement the limitations of LLMs, particularly in tasks with vast search spaces. Conversely, LLMs contribute with rich prior knowledge, guiding searches and directly handling text, addressing some shortcomings of EAs in text-related tasks. We introduced three research paradigms: LLM-enhanced Evolutionary Optimization, EA-enhanced LLM, and integrated synergy of LLM and EA. These paradigms exemplify the amalgamation of LLMs and EAs in various application scenarios, showcasing the potential collaborative strengths in tasks such as neural architecture search, code generation, software engineering, and text generation. As we look forward, the collaboration between EA and LLM has garnered increasing attention, with research focusing on enhancing mutual performance and driving task-specific improvements. Our contributions include a systematic survey of the cross-disciplinary research between LLMs and EAs, categorization into three research paradigms, and a critical analysis of existing methods, identifying challenges and inspiring future research in this dynamic domain. This paper serves as the first comprehensive and up-to-date review specifically focused on the EA research in the Era of LLMs, providing insights into the potential synergies between these two powerful paradigms for tackling complex optimization problems in the evolving landscape of artificial intelligence.

### v3

@@ IV-D Other Generative Tasks
IV-D1 Text Generation
Text generation is one of the most direct applications of LLMs. Xiao et al. used the generation ability of LLM and the search ability of EA for news summary generation [25], where event patterns structure is employed to represent and describe the key information, relationships and characteristics of events. They used LLM to extract event patterns from text, and then used GA to evolve the event pattern pool according to the importance and information quantity of argument roles, selecting the event pattern with the highest fitness to input into LLM for news summary generation. Some studies draw inspiration from text generation to optimize non-textual domains. Lim et al. developed the SCAPE system to explore conceptual architecture design by combining EA and LLM [147]. SCAPE represents design schemes or concepts using text, which allows EA to endow LLM with stronger innovative capabilities. The system enables architects to participate in parent selection, significantly improving the quality and exploration efficiency compared to vanilla generative AI tools. Another interesting application is to generate Super Mario game levels using LLMs [148, 21]. Sudhakaran et al. used a string representation method similar to the Video Game Level Corpus dataset [149] to encode levels, and then trained a LLM namely MarioGPT based on GPT-2 that can generate Mario levels according to natural language descriptions. By combining MarioGPT with the Novelty Search algorithm [150], an open-ended level generation process was formed. Novelty search can continually discover levels with diverse structures and gameplay styles.
IV-D2 Text-to-image Generation
In this type of task, when using LLMs like Stable Diffusion [151] to generate photorealistic images, it is necessary to optimize the input text prompts and model parameters in order to achieve better performance. Berger et al. tried to solve this problem and proposed the StableYolo method [152]. Its main innovation lies in combining EAs with computer vision for the first time, simultaneously optimizing language prompts and model hyperparameters to improve image generation quality. Among them, the LLM Stable Diffusion plays the main role of image generation, generating a series of images based on the input text prompts. While the GA plays the role of searching and optimizing. It uses the recognition confidence given by the YOLO object detection model as the fitness function to perform multi-objective optimization. The optimization objects include keywords in text prompts and various hyperparameters of the Stable Diffusion model. Through repeated iteration, the best language prompts and model settings that maximize image quality are found.
In addition to text-related tasks, some applications in natural sciences and social sciences are provided as follows to demonstrate the combined advantages of LLMs and EAs.
IV-D3 Natural Science
@@
IV-D4 Social Science
Suzuki et al. applied LLMs and EAs to study the evolutionary dynamics of social populations under conflicts of interest and behaviors between individuals [156], i.e. the propagation and changes of different personality traits and behavioral strategies. They used LLMs to make trait expressions more complex and higher-order, able to directly describe personality and psychological attributes that are difficult to model directly, and map them to behaviors. They then constructed an agent-based model where each agent uses a short natural language description to represent a personality trait related to cooperation as its gene. The LLM was used to derive a behavioral strategy for each agent based on the personality trait description and game history. The population evolved according to a GA, with offspring inheriting mutated versions of their parents’ trait genes, thus simulating the evolutionary dynamics of population changes over generations. This study provided a novel approach for researching the evolution of complex social populations.
In summary, the collaboration between LLMs and EAs has driven advancements in distinct generative task domains. By exploiting their strengths, researchers have made significant progress in these domains, paving the way for further advancements in generative modeling and related fields.
@@ V Roadmap and Future Directions
In the previous sections, we have reviewed the recent advances in unifying LLMs and EAs, but there are still many challenges and open problems that need to be addressed. In this section, we discuss the future directions of this research area. As depicted in Fig. 3, our perspective for future research includes investigations into methods, theory, and applications. Some of these studies aim to build upon the strengths and weaknesses highlighted in the reviewed research to propose more advanced approaches. Meanwhile, other research endeavors seek to explore new frontiers that go beyond the current body of knowledge, aiming to forge new research questions and infuse the field with renewed energy. This section will provide a separate introduction to these future research directions.
Figure 3: Roadmap and Future Directions.
V-A LLM-assisted EA for Complex & Large-scale Optimization
@@
Existing research has partially validated the ability of LLMs to address small-scale numerical optimization problems [29, 30, 22], but practicality still poses challenges. Firstly, for complex optimization problems with high-dimensional search spaces, numerous constraints, and high-precision numerical optimization, the limited context understanding and prompt length restrictions of large models may increase the difficulty of interaction with LLMs. Secondly, as black-box optimizers, the decision-making process of LLMs is difficult to interpret [157], and it remains unknown whether their optimization ability stems from reasoning or randomness. Moreover, the scope of the evaluated problems in existing research is relatively narrow [29, 30, 22], often limited to specific optimization problems, and the assessments only consider limited influencing factors, which are insufficient to fully demonstrate the optimization capabilities of LLMs. Another issue is that empirical studies suggest that LLMs may struggle to handle constrained problems effectively, making it crucial to incorporate constraint conditions into LLM optimization frameworks [28].
@@
Therefore, to further explore and utilize the capabilities of LLMs in solving optimization problems, future work can focus on addressing the aforementioned challenges and improving various aspects such as LLM behavior interpretation and evaluation, rational utilization of optimization information, and solving complex problems. Currently, efforts can be made to interpret and understand the behavior of LLMs, such as analyzing their internal attention mechanisms [158] or investigating the relationship between LLM temperature settings and exploration of optimal solutions, in order to gain a more objective understanding of LLM optimization abilities. Based on these foundations, improvement strategies or new pre-training models can be proposed to solve optimization problems with higher dimensions and complex constraint conditions. Beyond these research directions, the contextual understanding and processing abilities of LLMs can be further exploited, for example, by analyzing historical information of the optimization process and considering more advanced prompt methods [159]. Conducting validation on large-scale, real-world problems going forward would also provide meaningful value.
@@ V-B LLM-assisted Automatous and Intelligent EA
LLM holds significant promise for broader application scenarios within evolutionary optimization. In the following, we introduce several prospective areas, which, despite being relatively underexplored at present, are expected to drive more automatous and intelligent EAs in black-box optimization across diverse domains.
First, multi-modal LLMs provide opportunities for cross-domain EAs. Some LLMs have learned the relationships between different modalities during pre-training, allowing them to be used for downstream tasks involving multiple modalities [160, 161], such as image captioning and text-to-image generation. The ability of LLMs to understand correspondences across modalities can help EAs achieve cross-modal crossover and mutation by providing cross-modal prompts or examples. Additionally, population construction and evaluation can be improved across modalities. Furthermore, as LLMs technologies continue to progress, evolutionary operators based on LLMs will also benefit and improve the performance of EAs. With increasing model scales and stronger generalization capabilities, evolutionary operators based on LLMs will be better able to simulate more complex evolutionary mechanisms precisely. This provides potential for tackling problems in complex search spaces. The emerging and rapidly advancing field of LLMs will also provide more choices for EAs. LLMs with different pre-training characteristics will continue to enrich and enhance the capabilities of evolutionary operators based on LLMs.
@@
V-C Robust and Stable Prompt Engineering
@@
Regarding the issue of initialization, future research might consider multi-source initialization, simultaneously utilizing LLM to automatically expand the size and quality of the initial population. For problems with a high search space complexity, it is necessary to design more efficient evolutionary strategies, such as introducing richer evolutionary operators, leveraging the strengths of different EAs, and employing adaptive evolution. Another promising direction worth exploring is the utilization of human feedback to expedite and optimize the prompt discovery process, which has been overlooked in current research. In addressing the constraint of LLM on prompt performance, apart from considering the use of more powerful LLM models or designing dedicated fine-tuning methods, an alternative research direction at a lower level is to extract internal representations of LLM and study their mechanisms for understanding prompts, thereby guiding the design and representation of prompt strategies more effectively. Stable prompt optimization is also a promising direction to mitigate performance fluctuations caused by randomness.
V-D Theoretical Study for Specific LLM+EA Methods
Although empirical studies have validated the effectiveness of combining LLM and EA on small-scale problems, the incentives of their interaction are not yet clear. This reminds us to explore the sources of mutual promotion between LLM and EA in theoretical research, and to analyze in detail their complementary advantages and existing issues in large-scale empirical studies, thereby further promoting further improvement. This research can delve into two aspects:
V-D1 Algorithm Analysis
Researchers can analyze specific algorithms that combine LLM and EA and explore their convergence, complexity, and other properties. For convergence analysis, researchers can verify whether the algorithm can converge to the optimal solution or local optimum during the iterative process. This may involve proving the optimization properties of the objective function or fitness function during the algorithm’s iterative process and analyzing the impact of algorithm parameter settings on convergence. Regarding complexity analysis, future work can analyze the time and space complexity of algorithms that combine LLM and EA. This includes evaluating the complexity of key steps in the algorithm, such as generating new individuals, selection, crossover, and mutation operations. The applicability of methods that combine LLM and EA on different types of problems is another perspective could be studied, which may involve analyzing the relationship between problem attributes, constraints, problem size, and algorithm performance, as well as studying the performance guarantees or theoretical limits of algorithms on specific problem types.
V-D2 Optimization Theory
In problem modeling and analysis, specific problems that combine LLM and EA can be modeled and analyzed from the perspective of optimization theory. This may involve defining and characterizing the objective function, constraints, and exploring the feasible solution space of the problem. The key point of research lies in the search strategy, where researchers can study how to select and design appropriate search strategies to improve the performance of LLM and EA combination. This may include comparing and analyzing the effects of different search strategies (such as selection, crossover, and mutation operations) and sensitivity analysis of search strategy parameters. Moreover, future work can analyze the complexity of these methods on specific problems, including analyzing the computational complexity category of problems (such as P, NP, NP-hard) and analyzing the approximate performance of algorithms to determine the theoretical limits of algorithms in solving specific problems.
V-E Generalized Code Generation for Complex Task
@@
LLM and EA jointly contribute to another field of advancement, i.e., code generation. Currently, a plethora of research has emerged in this domain, which has further spurred the development of various downstream tasks, such as software engineering and EA design.
@@
V-F Generalized or Universal NAS
In the realm propelled by the collaboration of LLM and EA, NAS stands out as a vital application scenario. The prior knowledge about network architecture and code generation capability of LLM can greatly assist EA in efficiently discovering optimal architectures. However, existing work still faces various challenges. Some of these challenges are common among NAS methods based on EAs, such as high time consumption [162]. Additionally, the integration of LLM introduces new issues. Firstly, it is important to note that the current state-of-the-art LLM models, while excelling in various tasks, are not specifically tailored for NAS. Their ability to tackle NAS tasks stems from the presence of network architecture-related information in the training data [123]. Consequently, different LLM models exhibit significant variations in performance in NAS tasks. Furthermore, when compared to mainstream NAS methods, LLM-based approaches still exhibit some gaps in terms of their application scope and generalization ability. Additionally, certain ablation experiments reveal that the direct utilization of LLM prompts yields suboptimal learning results [125], underscoring the limitations of the current capabilities of LLM. Exploring fine-tuning methods that combine LLM with EAs has proven necessary.
V-G Applications and Innovations in Higher-level Tasks
In the discussed techniques within this paper, we observe their role as submodules in higher-level tasks. For instance, prompt optimization plays a role in LLM-based tasks, while code generation serves a purpose in NAS or software engineering tasks. In fact, these techniques can be utility in a broader range of higher-level tasks. Taking software engineering as an example, in higher-level tasks (such as repository-level software engineering), LLM and EA can fulfill more applicable demands. LLM can be used to implement cross-project collaboration. LLM not only can analyze the source codes and histories of multiple related projects to provide a large-scale search space for EA, but also its inherent prior knowledge can be used to identify the correspondence between knowledge of different projects, where EA can then utilize this relationship to perform knowledge transfer. In addition, EA can automatically generate version control strategies for different development stages, such as how to merge branches and when to release versions, balancing development efficiency and stability. Or EA can optimize multi-objective quality prediction models for different stages of the software lifecycle. In addition to software engineering, LLM and EA have the potential to promote the development of automatic circuit design and provide new innovative ideas for circuit designers. LLM can analyze a large amount of existing circuit design documents to extract key information such as functional modules and interface specifications to define the design space for EA. EA can optimize circuit topology, component selection and layout to meet multi-objective optimization of performance, power consumption and cost. LLM can also generate initial circuit design schemes for EA with reference to existing design cases. EA further optimizes the circuit topology through LLM-guided variation and selection. Among them, the design schemes and reports generated by LLM can be used as training data to improve LLM’s ability in this field. Moreover, with the development of microelectronics technology, the scale of circuits continues to expand, and LLM in large-scale input processing and EA in layered optimization may play an important role. Another example unmentioned before is the intelligent agent [163], which can learn and evolve its behavioral strategies through interaction with the environment. In [164], the inference and reflection capabilities of the intelligent agent are realized through interaction with LLM. Among them, policy optimization is achieved through prompt optimization, following a simplified evolutionary process similar to “natural selection”. Future research can explore more complex evolutionary mechanisms to aid in the self-evolution of intelligent agents.
@@ VI Conclusion
In this paper, we have undertaken a comprehensive exploration of the intersection between Evolutionary Algorithms (EAs) and Large Language Models (LLMs) in the transformative era of AI. We introduced three research paradigms: LLM-enhanced EA, EA-enhanced LLM, and applications driven by the integrated synergy of LLM and EA. These paradigms exemplify the amalgamation of LLMs and EAs in various application scenarios, showcasing their collaborative strengths in tasks such as Neural Architecture Search (NAS), code generation, software engineering, and text generation. As we look forward, the collaboration between EA and LLM has garnered increasing attention, with research focusing on enhancing mutual performance and driving task-specific improvements. Currently, existing LLMs lack the direct capability to handle complex and large-scale optimization problems, but they have shown promise in generating optimization algorithms and evaluating optimization results. Future work should focus on leveraging LLMs to solve more complex optimization problems by exploring approaches such as interpreting LLM behavior and comprehensively evaluating performance on diverse issues. The exploration of enhancing LLMs with EAs is still in its early and fragmented stages. Areas beyond prompt engineering and NAS deserve further in-depth and systematic investigation. These endeavors are crucial in advancing Evolutionary Computing to effectively address the forefront challenges in the AI community. Theoretical analysis of algorithms combining LLMs and EAs from an optimization theory perspective is essential to understand properties like convergence and complexity. In practical applications, mature collaborative paradigms have emerged where LLMs and EAs work together. These paradigms leverage the iterative search framework provided by EAs and the intelligent evolutionary operators offered by LLMs to tackle various text-based optimization problems, such as code generation and prompt engineering. Generalized code generation for intricate tasks may involve modular and interactive methods. Broadening the application scope and generalization ability of LLMs in NAS through techniques such as pretraining specialized models also merits investigation. While improvements within this established paradigm or extensions to new application domains are valuable, it is of paramount importance to transcend this paradigm and explore novel collaborative mechanisms. Overall, the collaboration between EA and LLM holds significant potential for future advancements. By further refining the interaction between these fields, we can pave the way for more efficient and effective optimization strategies and drive innovation in the broader field of artificial intelligence.